OpenStudy (anonymous):

What is x2 - 7x = -12 ?

OpenStudy (rose4825):

x2? do you mean: \[x^{2}\]

OpenStudy (anonymous):

X²-7x+12=0

OpenStudy (anonymous):

Can u continue with the hint sent to u above this message

OpenStudy (anonymous):

start with \[x^2-7x+12=0\] then factor

OpenStudy (anonymous):

How do I factor?

OpenStudy (anonymous):

What two values multiply to a positive 12 and sum to a negative 7?

OpenStudy (anonymous):

if 4X3 is 12 how can u use 4 and 3 to get -7

OpenStudy (anonymous):

-4 - 3?

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

so what is your answer

alones (alones):

Let's solve your equation step-by-step. x2−7x=−12 Step 1: Subtract -12 from both sides. x2−7x−(−12)=−12−(−12) x2−7x+12=0 Step 2: Factor left side of equation. (x−3)(x−4)=0 Step 3: Set factors equal to 0. x−3=0 or x−4=0 x=3 or x=4 I hope MY STEPS WILL HELP HERE!!